find 2 solutions for "n"
14n^2=25n+25
which leads to.. 14n^2-25n-25=0
help pleasee

Answer by scott8148(6628)

factoring __ (7n+5)(2n-5)=0

7n+5=0 __ 7n=-5 __ n=-5/7

2n-5=0 __ 2n=5 __ n=5/2